OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: 3 • (x • y) = (3 • x) • y Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

OpenStudy (zzr0ck3r):

(ab)c=a(cb) is associative property

OpenStudy (anonymous):

yea what he said

OpenStudy (anonymous):

zzrock3r thank you both

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

i got it right thankyough both again

OpenStudy (zzr0ck3r):

ab = ba is commutative a(b+c) = ab+ac is distributive a=b and b=c then a = c transitive

OpenStudy (anonymous):

thanks very much im closing it now have a good day.